Your answer is partially correct. Try again. The outstanding capital stock of Indigo Corporation consists of 1,900 shares of $100 par value, 9% preferred, and 5,400 shares of $50 par value common. Assuming that the company has retained earnings of $87,500, all of which is to be paid out in dividends, and that preferred dividends were not paid during the 2 years preceding the current year, state how much each class of stock should receive under each of the following conditions. (a) The preferred stock is noncumulative and nonparticipating. (Round answers to o decimal places, e.g. $38,487.) Preferred Common 17100 70400 (b) The preferred stock is cumulative and nonparticipating. (Round answers to O decimal places, e.g. $38,487.) Preferred Common 51300 36200 (c) The preferred stock is cumulative and participating. (Round the rate of participation to 4 decimal places, e.g. 1.4278%. Round answers to 0 decimal places, e.g. $38,487.) Preferred Common 4915 6985
Answer #1

Ans:

1900 *100 *9%=190,000*9%=17100

5400 shares 50 par value=$270,000

Preferred dividend

Common Dividend

Case a

17100

87,500 - 17100 = $ 70,400

Case b

17100 for current year + (17,100 *2 ) for past 2 years

87,500 – 51,300

17100+ 34200 = $ 51,300

$ 36,200

Case c

56,215.22    [51,300+$4915.22]

31,284.78       [$ 24,300+$ 6984.78]

**   In case of participative preference shares , common stock dividend = rate of preferred dividend and then any surplus profit is distributed in ratio of capital investment

Computation of share of income under Case C:

Preferred stock = 1900 * 100 = $ 190,000

Common stock = 5400*50 =$ 270,000

Distribution of current dividend and past dividend to preference shares = 51,300

      Distribution of dividend to common stock = $ 270,000 * .09 =$ 24,300

surplus profit after distribution of dividend to both shareholders

     87,500 – 51,300 -24,300 = $ 11,900 in 190,000: 270,000

Preferred dividend = 11900 *190000 /(190000+270000)

                                       = 11900*190,000/ 460,000

                                         = $4915.22

common stock = 11,900 - $4915.22 = $ 6984.78